OpenStudy (ijustneedsomehelppleazzzzzzzzz):

A square pool has an area of 292 square feet. To the nearest tenth, what is the length of one side of the pool? A. 8.5 feet B. 17.1 feet C. 73.0 feet D. 146.0 feet

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

is it b

OpenStudy (_malice_):

I think it's C.

OpenStudy (ijustneedsomehelppleazzzzzzzzz):

ok

OpenStudy (_malice_):

I "think", don't trust my answers, just giving my input ;)

OpenStudy (_malice_):

Was I right?

OpenStudy (whpalmer4):

you have a square pool with area 292 square feet. The formula for the area of a square with sides of length \(s\) is \[A=s^2\] So you need to find a value of \(s\) such that \[s^2=s*s=292\] You have a number of choices here, and one of them is correct. It should not be difficult to figure which one it is. You can narrow it down quickly by trying some numbers which are easy to multiply. \[10*10=100\]not big enough \[15*15=225\]not big enough \[30*30=900\]too big